Hi,

On 3-10-2011 19:28, Stephen A. Lawrence wrote:


On 11-10-03 12:38 PM, Man on Bridges wrote:
Hi,

On 2-10-2011 14:25, Horace Heffner wrote:
In other words, if there is any possibility of failure, i.e. q<1, then repeated events eventually, much more quickly than ordinary common sense dictates, result in failure. For example, if the probability of a catastrophe when drinking and driving once is 1/1000 = .001, then the probability of *no* catastrophe in 100 such events is 1-.999^100 = 0.095, or about 1%.

This should of course be read as:
then the probability of *a* catastrophe in 100 such events is 1-.999^100 = 0.095, or about 1%.


Both wrong.  Where'd you guys get those 1% values, anyway?

1 - 0.001 = 0.999, sure enough.

But 0.999 ^ 100 = 0.904792147...., which means there's about 90% chance of no catastrophe, or about 10% chance of a mess. Strangely, neither of you got the 10% number, even though it's not just correct, it's also what you'd guess if you didn't know anything (1 chance in 1000, repeated 100 times, give about 1 chance in 10 of hitting).

Ah, I see I also misread Horrace's number 0.09520...; you are right about the 0.095 or 9.5%; Horrace and I both overlooked the position of the decimal point, but my point was that Horrace reversed q(success) and p(failure) values and if you add it to 0.90479... (success), you will see that the total is 1 or 100%. So the numbers are still right!
But that's why peer-reviews are important for those kind of matters.

Kind regards,

MoB

Reply via email to